# Cubic fourfolds with symplectic automorphisms

### v1

@@ 6.6. The group $\mathbf{Hol}(5)$
[…]
$\begin{array}[]{c|rrrrr}\rm class&\rm 1&(12)(34)&(2345)&(2534)&(12354)\cr\rm size&1&5&5&5&4\cr\hline\cr\chi_{1}&1&1&1&1&1\cr\chi_{2}&1&1&-1&-1&1\cr\chi_{3}&1&-1&-i&i&1\cr\chi_{4}&1&-1&i&-i&1\cr\chi_{5}&4&0&0&0&-1\cr\end{array}$
@@ 7.2. The projective special linear group $\mathrm{L}_{2}(7)$
in $\mathbb{P}^{6}$ where $x_{ijk}=x_{i}x_{j}x_{k}$. This family contains the Clebsch-Segre cubic (2.1), and a general menber is smooth.
@@ 7.3. The symmetric group $\mathrm{S}_{5}$
and this satisfies the symplectic conditions. We have $\pi_{1}|\mathrm{A}_{5}=\rho_{1}$ and $V_{3}(\rho_{1})$ is decomposed into (±$(\pm 1)$-eigenspaces for $(12)\in\mathrm{S}_{5}$
@@
A general menber of $V_{3}(\pi_{1})$ defines a smooth cubic 4-fold since $y^{2}s_{1}(x)+s_{3}(x)=0$ is smooth, but $F\in V_{3}(\pi_{1}\otimes\eta_{2})$ is reducible. We have $\dim\mathrm{C}_{\pi_{1}}(\mathrm{S}_{5})=1^{2}+1^{2}+1^{2}=3$ and
@@
Next we consider a representation $\pi_{2}$ affording $\phi_{2}$. Let us consider the following involution with eigenvalues $(1,1,1,1,-1,-1)$:
@@
A general menber of $V_{3}(\pi_{2})$ defines a smooth cubic 4-fold since $s_{1}(x)s_{2}(x)+(I_{1}-I_{2})=0$ is smooth. On the other hand,
@@
is singular at $[1:\cdots:1]$. We have $\dim\mathrm{C}_{\pi_{2}}(\mathrm{S}_{5})=1^{2}+1^{2}=2$ and
@@ 7.4. The Mathieu group $\mathrm{M}_{9}$
We have $\mathrm{M}_{9}\cong 3^{2}.\mathrm{Q}_{8}$ and this group contains $43^{2}.$4 in section 6.4 as a subgroup of index 2. The Schur multiplier of $\mathrm{M}_{9}$ is $\mathrm{C}_{3}$, and we must consider a Schur cover $3.\mathrm{M}_{9}$. The character table of $\mathrm{M}_{9}$ itself is as follows:
@@
[…]
@@ 7.5. The group $\mathbf{N_{72}\cong(S_{3}\times S_{3}):2}$ with GAP-ID (72,40)
$\begin{array}[]{c|rrrrrrrrr}\rm class&\rm 1&\rm 2A&\rm 2B&\rm 2C&\rm 3A&\rm 3B&\rm 4A&\rm 6A&\rm 6B\cr\rm size&1&6&6&9&4&4&18&12&12\cr\hline\cr\xi_{1}&1&1&1&1&1&1&1&1&1\cr\xi_{2}&1&-1&-1&1&1&1&1&-1&-1\cr\xi_{3}&1&-1&1&1&1&1&-1&-1&1\cr\xi_{4}&1&1&-1&1&1&1&-1&1&-1\cr\xi_{5}&2&0&0&-2&2&2&0&0&0\cr\xi_{6}&4&-2&0&0&1&-2&0&1&0\cr\xi_{7}&4&0&-2&0&-2&1&0&0&1\cr\xi_{8}&4&0&2&0&-2&1&0&0&-1\cr\xi_{9}&4&2&0&0&1&-2&0&-1&0\cr\end{array}\qquad\begin{array}[]{l}\rm 2A=(12)\\ \rm 2B=(14)(25)(36)\\ \rm 2C=(12)(45)\\ \rm 3A=(123)\\ \rm 3B=(123)(456)\\ \rm 4A=(14)(2536)\\ \rm 6A=(12)(456)\\ \rm 6B=(142536)\\ \end{array}$
@@
A representation $\pi$ affording the character $\psi$ is given by extending $\rho_{1}$ in secton 6.5. Let us consider a matrix $c=\begin{bmatrix}C_{1}&C_{2}\\ C_{2}&C_{1}\end{bmatrix}$ of order 2 with eigenvalues $(1,1,1,1,-1,-1)$, where
@@ 7.6. The group $\mathbf{T_{48}\cong Q_{8}:S_{3}}$ with GAP-ID (48,29)
of order 2 with eigenvalues $(1,1,1,1,-1,-1)$. They normalize $\rho(\mathrm{Q}_{8})$, and $n_{1},n_{2}$ together with $\rho(\mathrm{Q}_{8})$ generate a group $H\cong\mathrm{Q}_{8}:\mathrm{S}_{3}$ with GAP-ID (48,29) (Conjugations by $n_{1}$ and $n_{2}$ permute conjugacy classes
@@
A general menber of $V_{3}(\pi)$ is smooth since sum of the above four generator define a smooth cubic fourfold. We have $\dim\mathrm{C}_{\pi}(\mathrm{T}_{48})=1^{2}+1^{2}+1^{2}=3$ and
@@ 8. Maximal groups
of order 4 with eigenvalues $(1,1,-1,-1,i,-i)$. The group $H\cong 3.\mathrm{A}_{6}$ in section 2.5 is normalized by $r$, and $H$ together with $r$ generate $3.\mathrm{M}_{10}$. By the action of $r$, the space $V_{3}(\rho)=\mathbb{C}f(x)\oplus\mathbb{C}g(x)$ is decomposed into (±)plus-or-minus($\pm$)( ± )-eigenspaces $V_{3}(\rho)_{\pm}=\mathbb{C}F_{\pm}(x)$ where

### v2

@@ 6.6. The group $\mathbf{Hol}(5)$
The group $\mathrm{Hol}(5)\cong\mathrm{D}_{10}.2$ is given as a subgroup of $\mathrm{S}_{5}$ generated by $(12)(34)$ and $(2,3,4,5)$. The Schur multiplier of $\mathrm{Hol}(5)$ is 1, and hence we have $\mathrm{Hol}(5)^{*}=\mathrm{Hol}(5)$. The character table of $\mathrm{Hol}(5)$ is as follows:
$\begin{array}[]{c|rrrrr}\rm class&\rm 1&(12)(34)&(2345)&(2534)&(12354)\cr\rm size&1&5\quad&5\quad&5\quad&4\quad\cr\hline\cr\chi_{1}&1&1\quad&1\quad&1\quad&1\quad\cr\chi_{2}&1&1\quad&-1\quad&-1\quad&1\quad\cr\chi_{3}&1&-1\quad&-i\quad&i\quad&1\quad\cr\chi_{4}&1&-1\quad&i\quad&-i\quad&1\quad\cr\chi_{5}&4&0\quad&0\quad&0\quad&-1\quad\cr\end{array}$
@@ 7.2. The projective special linear group $\mathrm{L}_{2}(7)$
in $\mathbb{P}^{6}$ where $x_{ijk}=x_{i}x_{j}x_{k}$. This family contains the Clebsch-Segre cubic (2.1), and a general member is smooth.
@@ 7.3. The symmetric group $\mathrm{S}_{5}$
and this satisfies the symplectic conditions. We have $\pi_{1}|\mathrm{A}_{5}=\rho_{1}$ and $V_{3}(\rho_{1})$ is decomposed into $(\pm 1)$-eigenspaces for $(12)\in\mathrm{S}_{5}$
@@
A general member of $V_{3}(\pi_{1})$ defines a smooth cubic 4-fold since $y^{2}s_{1}(x)+s_{3}(x)=0$ is smooth, but $F\in V_{3}(\pi_{1}\otimes\eta_{2})$ is reducible. We have $\dim\mathrm{C}_{\pi_{1}}(\mathrm{S}_{5})=1^{2}+1^{2}+1^{2}=3$ and
@@
Next we consider a representation $\pi_{2}$ affording $\phi_{2}$. Let us consider the following involution with eigenvalues (1,1,1,1,1,1)(1,1,1,1,-1,-1):
@@
A general member of $V_{3}(\pi_{2})$ defines a smooth cubic 4-fold since $s_{1}(x)s_{2}(x)+(I_{1}-I_{2})=0$ is smooth. On the other hand,
@@
is singular at $1:\cdots:1]$. We have $\dim\mathrm{C}_{\pi_{2}}(\mathrm{S}_{5})=1^{2}+1^{2}=2$ and
@@ 7.4. The Mathieu group $\mathrm{M}_{9}$
We have $\mathrm{M}_{9}\cong 3^{2}.\mathrm{Q}_{8}$ and this group contains $3^{2}.4$ in section 6.4 as a subgroup of index 2. The Schur multiplier of $\mathrm{M}_{9}$ is $\mathrm{C}_{3}$, and we must consider a Schur cover $3.\mathrm{M}_{9}$. The character table of $\mathrm{M}_{9}$ itself is as follows:
@@
are satisfied only by $\chi_{15}$ and $\chi_{16}=\overline{\chi_{15}}$. Representaions affording $\chi_{15}$ and $\overline{\chi_{15}}$ are given by extending the action of $H_{2}\cong 3.(3^{2}.4)$ in section 6.4. Let us consider a matrix $g=\begin{bmatrix}{\bf 0}&h_{2}\\ \omega h_{2}&{\bf 0}\end{bmatrix}$ of order 4 with eigenvalues (1,1,1,1,i,i)(1,1,-1,-1,i,-i) where
@@ 7.5. The group $\mathbf{N_{72}\cong(S_{3}\times S_{3}):2}$ with GAP-ID (72,40)
[…]
@@
A representation $\pi$ affording the character $\psi$ is given by extending $\rho_{1}$ in secton 6.5. Let us consider a matrix $c=\begin{bmatrix}C_{1}&C_{2}\\ C_{2}&C_{1}\end{bmatrix}$ of order 2 with eigenvalues (1,1,1,1,1,1)(1,1,1,1,-1,-1), where
@@ 7.6. The group $\mathbf{T_{48}\cong Q_{8}:S_{3}}$ with GAP-ID (48,29)
of order 2 with eigenvalues (1,1,1,1,1,1)(1,1,1,1,-1,-1). They normalize $\rho(\mathrm{Q}_{8})$, and $n_{1},n_{2}$ together with $\rho(\mathrm{Q}_{8})$ generate a group $H\cong\mathrm{Q}_{8}:\mathrm{S}_{3}$ with GAP-ID (48,29) (Conjugations by $n_{1}$ and $n_{2}$ permute conjugacy classes
@@
A general member of $V_{3}(\pi)$ is smooth since sum of the above four generator define a smooth cubic fourfold. We have $\dim\mathrm{C}_{\pi}(\mathrm{T}_{48})=1^{2}+1^{2}+1^{2}=3$ and
@@ 8. Maximal groups
of order 4 with eigenvalues (1,1,1,1,i,i)(1,1,-1,-1,i,-i). The group $H\cong 3.\mathrm{A}_{6}$ in section 2.5 is normalized by $r$, and $H$ together with $r$ generate $3.\mathrm{M}_{10}$. By the action of $r$, the space $V_{3}(\rho)=\mathbb{C}f(x)\oplus\mathbb{C}g(x)$ is decomposed into $(\pm)$-eigenspaces $V_{3}(\rho)_{\pm}=\mathbb{C}F_{\pm}(x)$ where
@@ Appendix A An outer automorphism of $\mathrm{S}_{6}$
Appendix B Table
For non-maximal cases of Laza and Zheng’s 34 groups, we have the following table, where $r(G)=\mathrm{rank}\ S_{G}$ (see [LZ22]), GAP-ID is “IdGroup(GG)” in GAP, $M$ is the Schur multiplier of $G$, and ID* is GAP-ID of a Schur covering group $G^{*}$ given by “SchurCover(GG)” in GAP. Equations of corresponding cubic 4-folds are given in Eq, and irr is the number of irreducible components of the moduli space $\mathcal{M}(G)$ of cubic fourfolds $X$ such that $G\subset\mathrm{Aut}^{s}(X)$.
Table 2. Schur covers and multipliers
$r(G)$ $G$ GAP-ID $M$ ID* Eq irr
8 2 2,1 1 2,1 (3.1) 1
12 $2^{2}$ 4,2 2 8,3 (3.2) 1
12 3 3,1 1 3,1 (3.3), (3.4) 2
14 4 4,1 1 4,1 (3.5) 1
14 $\mathrm{S}_{3}$ 6,1 1 6,1 (3.6) 1
15 $\mathrm{D}_{8}$ 8,3 2 16,7 (3.7) 1
16 $\mathrm{A}_{3,3}$ 18,4 3 54,8 (4.6), (4.11) 2
16 $\mathrm{D}_{12}$ 12,4 2 24,8 (4.2), (4.3) 2
16 $\mathrm{A}_{4}$ 12,3 2 24,3 (4.20), (4.22) 2
16 $\mathrm{D}_{10}$ 10, 1 1 10, 1 (4.1) 1
17 $\mathrm{S}_{4}$ 24,12 2 48,28 (5.2), (5.3) 1
17 $\mathrm{Q}_{8}$ 8,4 1 8,4 (5.1) 1
18 $3^{1+4}_{+}:2$ 486, 249 $3^{5}$ - (6.1) 1
18 $\mathrm{A}_{4,3}$ 72,43 6 432,258 (6.5), (6.7) 2
18 $\mathrm{A}_{5}$ 60, 5 2 120,5 (6.9), (6.10) 2
18 $3^{2}.4$ 36,9 3 108,15 (6.11), (6.12) 2
18 $\mathrm{S}_{3,3}$ 36,10 2 72,22 (6.13) 1
18 $\mathrm{F}_{21}$ 21,1 1 21,1 (6.2) 1
18 $\mathrm{Hol}(5)$ 20, 3 1 20,3 (6.15) 1
18 $\mathrm{QD}_{16}$ 16,8 1 16,8 (6.3) 1
19 $3^{1+4}_{+}:2.2$ 972, 776 $3^{4}$ - (7.1) 1
19 $\mathrm{A}_{6}$ 360,118 6 - (2.6), (2.9) 2
19 $\mathrm{L}_{2}(7)$ 168,42 2 336,114 (7.2) 1
19 $\mathrm{S}_{5}$ 120,34 2 240, 89 (7.3), (7.6) 2
19 $\mathrm{M}_{9}$ 72,41 3 216,88 (7.9) 1
19 $\mathrm{N}_{72}$ 72,40 2 144,117 (7.10) 1
19 $\mathrm{T}_{48}$ 48,29 1 48,29 (7.12) 1